# =COLI-manual3.qry Cross-lingual Cross-temporal Summarization: Dataset, Models, Evaluation

### v1

@@ 5.1 Automatic evaluation metrics
Unsupervised extractive pipelines
Finding 1. Slightly Diminishing results from historical text normalization: We observe a decrease in all metrics scores on MemSum pipeline normalized with Norma for both hDe-En and hEn-De. This is contrary to our hypothesis that spelling normalization of historical texts can boost the translation quality for translators originally trained for modern text. Even though Norma performs well at token-level normalization, it fails to consider the context of the words, where in our use case some historical spellings may correspond to multiple modern meanings depending on the context. An example is the word “in” which can function as an adverb in the phrase e.g., “in den Wald” (in the forest) or the accusative case of the word “ihn” (him) in historical texts. A false normalization of such cases results in the meaning change of the entire sentence.
Finding 2. Positive gain from the retrieve-then-summarize pipeline with ChatGPT:
Compared to the extract-then-translate
pipeline, we observe an increase in metric scores for both hDe-En (all metrics except for BARTScore) and hEn-De (except for ROUGE-1). For hDe-En, it is scored as one of the top two models by ROUGE-L, BERTScore
and MENLI variants (in bold). This pipeline is competitive as it not only performs better but expands ChatGPT’s summarization potential for long documents.
Supervised abstractive e2e Finding 1. Mixed results on intermediate task finetuning: For hDe-En, intermediate task finetuning with external sources does not improve scores, and sometimes finetuning with more tasks even deteriorates the outcome according to the metrics. This may be because our CLCTS hDe-En dataset, which dates back to the early 1800s, has substantial linguistic differences to the external sources as discussed in Section 3. For hEn-De, the results show the opposite. We observe higher scores for models trained with more intermediate summarization tasks from external sources. Though, according to MENLI variants and BARTScore, it is not always better to add CTS tasks. This may be because the CTS tasks from HISTSUMM contains documents which have up to 200 years of difference in publication time and again lead to substantial linguistic differences.
@@
For hDe-En, the best finetuned model is mLED-Historical MT, where we finetune intermediately with our historical translation task before our target CLCTS task, though according to the metrics, the improvement is very small. Similarly, for the other direction hEn-De, the historical translation task improve model scores but only mildly compared to intermediate tasks. Since the historical translation dataset contains only 201 pairs, a larger size of the historical translation dataset may further improve the results.
Zero-shot abstractive ChatGPT
Finding 1. A close match between pipeline and e2e prompts: For both directions, we observe slightly better scores for summaries generated with e2e prompts which tell ChatGPT to summarize in the target language directly (see Table 6). However, in the following Section 6,
we show that the e2e prompts are prone to invalid outputs.
Finding 2. ChatGPT can output summaries from its memory: For the title prompt experiment, we input zero text from the document but only the title, author and publication year. Slightly surprisingly, for both directions, the resulting scores are mostly comparable to that of the supervised extractive models if not better. ChatGPT is capable of outputting summaries based solely on meta-information for all of our documents. This raises our question on the actual summarization ability of ChatGPT and therefore in the following Section 6,
@@
Comparison among all approaches
All the metrics give lower scores to extract-then-translate models (with/without normalization) and ChatGPT title prompt than other models, except for BARTScore, which rates the MemSum pipeline as the highest for hDE-En. However, mixed results are observed from the metrics for determining the best models. BERTScore variants in both directions give higher scores to ChatGPT-related models (excluding title prompt). MENLI variants agree with this fully for hDe-En, and partially agree for hEn-De. The same is also observed from MoverScore and ROUGE-L. BARTScore, however, contradicts with them and scores ChatGPT models higher for hEn-De, though only partly for hDe-En. Moreover, for hEn-De, ChatGPT is no longer dominating. However, the study from Goyal, Li, and Durrett (2022) also shows that current metrics fail to evaluate zero-shot model output properly. To better understand the performance of models and the metrics, we also consider human evaluation.
@@ 5.2 Human and ChatGPT evaluation
-title prompt | 2.88/2.75 | 3.25/2.25 | 4.75/3.25 | 2.31/2.25 |
-e2e prompt | 3.80/3.79 | 3.50/4.23 | 4.00/3.87 | 4.00/4.08 |
-pipeline prompt | 4.50/4.25 | 4.38/4.50 | 5.00/4.00 | 4.25/4.25 |
@@
-title prompt | 3.12/2.50 | 2.12/2.25 | 2.81/3.00 | 2.12/2.00 |
-e2e prompt | 3.75/3.21 | 3.13/3.18 | 3.25/3.36 | 3.13/2.85 |
-pipeline prompt | 3.06/3.25 | 3.38/3.25 | 3.06/3.50 | 3.38/3.25 |
@@
We report the mean value of human and ChatGPT annotations for each model in Table 9. The best scores for supervised abstractive models in both directions range between 2 to 3 by human annotators and only above 1 by ChatGPT, which indicates the quality of outputs is bad to mediocre. However, the highest scores for ChatGPT as a summarizer are over 4 (very good) for hDe-En and 3 to 4 (mediocre to good) for hEn-De. Overall, the quality of outputs for hEn-De is lower than that for hDe-En. This is also expected since the length of historical English documents is much longer with 9643.2 words on average which results in more information loss due to text truncation because of computational limitations and the cross-lingual embedding distance from Table 3 also suggests more difficulty of hEn-De compared to hDe-En.
@@
For both directions, humans give higher scores to summaries generated by ChatGPT, especially ChatGPT pipeline prompt for hDe-En with an increase of a great margin. Unlike evaluation metrics, humans prefer pipeline prompt output better than e2e prompt output for hDe-En and for hEn-De in coherence and relevance. A similar finding also holds for comparing finetuned models, where humans give higher scores for the model finetuned with historical translation similar to evaluation metrics for hDe-En. The ratings for hEn-De differs slightly
where human ratings agree with metric evaluation results in terms of coherence and consistency. Since the overall quality of generated summaries is bad for hEn-De, the comparisons among outputs may be more difficult for both humans and evaluation metrics.
Mixed results for ChatGPT as an evaluator.
@@
Table 9 (c) shows the annotation agreement between humans and ChatGPT annotation. For hDe-En, the agreement is still moderate but weak for hEn-De overall. Similarly to the findings from Chiang and yi Lee (2023), we observe lower Likert scores from ChatGPT evaluation compared to human evaluation except for summaries generated by ChatGPT itself. However, the output for hEn-De is more concerning where we observe many ratings of score 1, especially for relevance. Another point worth noticing is that even though in our instruction, we allow for 0.5 increments, we do not observe such ratings from ChatGPT output.
@@
To sum up, ChatGPT as an evaluator (i) is prone to lower scores than humans 101010We include one sample annotation from the gold standard annotations during instruction.; (ii) favours its own output since we observe similar (even higher) scores from ChatGPT evaluation compared to human evaluation for ChatGPT summarization outputs; (iii) cannot distinguish summaries of low qualities (hEn-De); (iv) does not fully follow the instruction of 0.5 score increments.

### v3

@@ 5.1 Automatic evaluation metrics
Supervised extractive pipelines
Finding 1. Slightly diminishing results from historical text normalization: We observe a decrease in all metrics scores on MemSum pipeline normalized with Norma for both hDe-En and hEn-De. This is contrary to our hypothesis that spelling normalization of historical texts can boost the translation quality for translators originally trained for modern text. Even though Norma performs well at token-level normalization, it fails to consider the context of the words, where in our use case some historical spellings may correspond to multiple modern meanings depending on the context. An example is the word “in” which can function as an adverb in the phrase for example, “in den Wald” (in the forest) or the accusative case of the word “ihn” (him) in historical texts. A false normalization of such cases results in the meaning change of the entire sentence.
Supervised abstractive e2e Finding 1. Mixed results on intermediate task finetuning: For hDe-En, intermediate task finetuning with external sources does not improve scores, and sometimes finetuning with more tasks even deteriorates the outcome according to the metrics. This may be because our CLCTS hDe-En dataset, which dates back to the early 1800s, has substantial linguistic differences from the external sources as discussed in Section 3. For hEn-De, the results show the opposite. We observe higher scores for models trained with more intermediate summarization tasks from external sources. Though, according to MENLI variants and BARTScore, it is not always better to add CTS tasks. This may be because the CTS tasks from HISTSUMM contain documents that have up to 200 years of difference in publication time and again lead to substantial linguistic differences.
@@
For hDe-En, one of the best finetuned models is mLED-Historical MT, where we finetune intermediately with our historical translation task before our target CLCTS task, though according to the metrics, the improvement is very small. Similarly, for the other direction hEn-De, the historical translation task improves model scores but only mildly compared to intermediate tasks. Since the historical translation dataset contains only 201 pairs, a larger size of the historical translation dataset may further improve the results.
Zero-shot abstractive ChatGPT Finding 1. A close match between pipeline and e2e prompts: For hDe-En, we observe slightly better scores for summaries generated with e2e prompts which tell ChatGPT to summarize in the target language directly (see Table 7). However, in the following Section 6, we show that the e2e prompts are prone to invalid outputs.
Finding 2. Competitive results from the retrieve-then-summarize design with ChatGPT: We observe competitive results from retrieve-then-summarize design for both hDe-En and hEn-De. Especially for hDe-En, it is scored as one of the top two models by ROUGE-L, MoverScore, BERTScore and MENLI. Also worth noting is that this design not only performs well but also expands ChatGPT’s potential for long document summarization.
Finding 3. ChatGPT can output summaries from its memory: For the e2e (Title) prompt experiment, we input zero text from the document but only the title, author, and publication year. Slightly surprisingly, for both directions, the resulting scores are mostly comparable to that of the supervised extractive models if not better. ChatGPT is capable of outputting summaries based solely on meta-information for all of our documents. This raises our question on the actual summarization ability of ChatGPT and therefore in the following Section 6,
@@
Comparison among all approaches All the metrics give lower scores to extract-then-translate models (with/without normalization) and ChatGPT e2e (Title) prompt than other models, except for BARTScore, which rates the MemSum pipeline as the highest for hDE-En. However, mixed results are observed from the metrics for determining the best models. BERTScore variants in both directions give higher scores to ChatGPT-related models (excluding e2e (Title) prompt). MENLI variants agree with this fully for hDe-En and partially agree for hEn-De. The same is observed from MoverScore and ROUGE-L. BARTScore contradicts with them and scores ChatGPT models higher for hEn-De, though not for hDe-En. Moreover, for hEn-De, ChatGPT is no longer dominating. However, the study from Goyal, Li, and Durrett (2022a) also shows that current metrics fail to evaluate the zero-shot model output properly. To better understand the performance of models and the metrics, we also consider human evaluation.
According to our automatic metrics, we find (1) extractive pipelines perform the worst among all methods; (2) supervised abstractive e2e methods outperform pipeline methods, where the historical translation task improves model performance mildly for both directions and we speculate that the effect of intermediate task finetuning depends on the linguistic similarity of external sources and target sources; (3) zero-shot abstractive ChatGPT delivers the best results under e2e prompts but can output summaries from its memory based on the experiment results from e2e (Title) prompts.
@@ 5.2 Human and ChatGPT evaluation
-e2e (Title) prompt | 2.55/1.70 | 3.05/1.20 | 4.10/3.70 | 2.02/1.40 |
-e2e prompt | 4.14/3.30 | 3.98/2.40 | 4.18/3.80 | 3.97/2.70 |
-pipeline prompt | 4.35/3.30 | 4.30/2.40 | 4.55/4.00 | 4.35/3.00 |
- | 4.08/2.90 | 4.08/2.10 | 4.15/3.50 | 4.25/2.70 |
@@
-e2e (Title) prompt | 3.31/2.00 | 2.25/1.00 | 3.28/3.50 | 2.31/1.38 |
-e2e prompt | 3.31/3.00 | 2.80/2.22 | 3.38/3.56 | 2.98/2.33 |
-pipeline prompt | 3.34/2.86 | 3.19/2.00 | 3.47/3.57 | 3.31/2.57 |
- | 3.38/3.00 | 2.94/2.38 | 3.34/3.75 | 3.06/2.62 |
@@
We report the mean value of human and ChatGPT (GPT-4-1106-preview) annotations for each model in Table 11.151515We utilize both GPT-3.5-turbo and GPT-4-1106-preview for annotation. GPT-4-1106-preview correlates better with human annotators than GPT-3.5-turbo. We report the annotation results from GPT-4-1106-preview in Table 11 in the main text. The annotation results from GPT-3.5-turbo are reported in Table 21 in Section 9.6 (appendix). The best scores for supervised abstractive models in both directions range between 2 to mildly above 3 by human annotators and below 2 by ChatGPT, which indicates the quality of outputs is bad to mediocre. However, the highest scores for ChatGPT as a summarizer are over 4 (very good) for hDe-En and 3 to 4 (mediocre to good) for hEn-De by human annotators. Overall, the quality of outputs for hEn-De is lower than that for hDe-En. This is also expected since the length of historical English documents is much longer with 9,643 words on average which results in more information loss due to text truncation because of computational limitations161616We also experimented with mBART finetuning with unlimiformer under both low-cost training and computationally expensive long-range training method. Since the low-cost training method only sees full texts at test time and long-range training suffers from the same GPU memory constraints as mLED, we do not observe improvements in model performance compared to baseline mLED. and the cross-lingual embedding distance from Table 3 also suggests more difficulty of hEn-De compared to hDe-En.
@@
For both directions, humans give higher scores to summaries generated by ChatGPT, especially ChatGPT pipeline prompt for hDe-En with an increase of annotation score to a great margin. Unlike the evaluation metrics, humans prefer pipeline prompt output over e2e prompt output for hDe-En and regarding consistency, fluency and relevance for hEn-De. For comparing finetuned models, humans give higher scores for the model finetuned with historical translation in coherence and consistency similar to evaluation metrics for hDe-En. The ratings for hEn-De differ where human ratings prefer mLED-Bidirection. Since the overall quality of generated summaries is worse for hEn-De, the comparisons among outputs may be more difficult for both humans and evaluation metrics.
@@
Mixed results for ChatGPT (GPT-4-1106-preview) as an evaluator. Table 21 (c) shows the annotation agreement between humans and ChatGPT annotation. For both directions, the agreement is moderate to good overall with a slightly better agreement for direction hEn-De (except for fluency). Similarly to the findings from Chiang and Lee (2023), we observe lower Likert scores from ChatGPT evaluation compared to human evaluation except for summaries generated by ChatGPT itself. However, the output for hEn-De is more concerning where we observe more ratings of score 1. Another point worth noticing is that even though in our instruction, we allow for 0.5 increments, we do not observe such ratings from ChatGPT output.
@@
To sum up, ChatGPT as an evaluator (1) is prone to lower scores than humans;171717We include one sample annotation from the gold standard annotations during instruction. (2) cannot distinguish summaries of low quality (hEn-De); (3) does not fully follow the instruction of 0.5 score increments.
5.3 Example outputs
[…]
For low quality generation (summary 1.2 and summary 2.2), we observe a substantial amount of factual inconsistency with the source text (1) and nonsensical text (3). In summary 2.2, we also notice the untranslated expression “Nimmersatt" repeatedly appears in the text. Unclear coreference of pronouns is another issue that affects the readability of the outputs. For example, in the sentence from summary 1.2 “the priest asks him if he (2) can come to the wedding”. The pronoun “he” is ambiguous as it could refer to the boy or the priest himself. In contrast, summary 1.1 and summary 2.1 contain fewer errors, especially the ChatGPT output which only contains one factual inconsistency. In summary 1.1, we find two factual inconsistencies with the source text where the boy “shares” the food instead of “cooking” it himself and he “asks” the image instead of “asking for” it. Additionally, we also notice several other errors such as unclear coreferences of pronouns, nonsensical texts, and spelling mistakes (such as “could’t”).
Reference Texts |
“A poor peasant boy hears a sermon that the way to heaven always leads straight ahead. He goes and comes to the service in a big city church, thinks he is in heaven and cannot be sent away. He sees people praying in front of Mary with the baby Jesus and thinks it is God. He shares his food with him, and the image grows thicker. Once, when he is bedridden for eight days, he goes and apologizes to the image. It invites him to the wedding next Sunday. The priest, who has been watching him, is not allowed to go, but gives him communion on Sunday, when the boy falls down dead.” |
1.1 Generated Text from model mLED-Base, quality level 3-4 |
“A poor peasant boy wants to go to Heaven, but the parson tells him to go straight. When he arrives in the church, he thinks God is there, and says, "Listen, dear God, I will bring you half of my food every day." Every day he cooks (1) half of his food, and the image begins to eat. A couple of weeks later, when people realize that the image is getting bigger and bigger, he stays in church and sees his bread being shared with the mother God. He gets sick and can’t get up for eight days. The parson goes after him and hears him say, "God, don’t take it so long (3), I was sick and could’t stand up." He answers (2), "That doesn’t matter, I’ve seen your good will, that’s enough, and next Sunday you are to come to my wedding." The boy rejoices and asks for the image (1). "No, it’s you alone," says he. He (2) first wants to prepare and gives him the supper. The next Sunday, when supper is coming to him, he falls down and is dead and is to eternal wedding." ” |
1.2 Generated Text from model mLED-MLS-CLS-CTS , quality level 1-2 |
“A poor peasant in the church hears God (1) admonishing him to go to Heaven, and he goes to God’s church. The priest tells him not to go out, but he wants to keep him (1). He stays in church and when he sees the mother God statue, which is cut from wood, the little Jesus (3), he says that it’s God, which means that he will eat half of his food every day (1). After a few weeks, people marvel at the statue and see how he spreads bread with God (1). After eight days, the priest asks him if he (2) can come to the wedding. The boy replies, "No, you’re only me, I told you that you told me that you were you alone. I’ll give him dinner first, which he says he will give him, and then he dies, then he’ll never be able to marry again," so he has to go back to the church and asks if he could come with God alone. After the wedding, the boy asks the parson to give him the dinner, but God says he’s dead and he is dead, and God says that he must marry him again (3).” |
Example 1 |
Reference Texts |
“The wolf is the stronger of the two and dominates the fox. The fox would like to be rid of him. The wolf forces the fox to get him food. He gets him a lamb and leaves. But the wolf is even hungrier and tries to get a lamb for himself. But he is so clumsy that the farmers notice and beat him. The wolf says to the fox that he has tricked him. The fox replied, ‘Why are you such a glutton!’ The next day the fox has to get pancakes for the wolf. They go to the house together and the fox sneaks around the house to carefully get six pancakes from the plate. The wolf again doesn’t have enough and wants to get more. But he again doesn’t manage and the farmer’s wife beats him again. The following day, the two go to the butcher. They jump through a window hatch into the cellar. The wolf eats a lot. The fox, on the other hand, eats little and keeps looking to see if he still fits through the window hatch. Suddenly the butcher comes, the red fox jumps out quickly, only the wolf gets stuck and is beaten to death by the butcher.” |
2.1 Generated Text from model ChatGPT-e2e prompt, quality level 4-5 |
“The wolf and the fox were walking in the woods together. The wolf demanded that the fox bring him food or he would eat him. The fox suggested they steal a lamb from a nearby farm. The wolf ate the lamb and wanted more, but when he tried to steal another, he was caught and beaten by the farmers. The next day, the fox suggested they steal pancakes from a farm, but the wolf ate too many and was caught again. On the third day, the fox suggested they steal salted meat from a cellar, but the wolf ate too much and got stuck in the cellar. The farmers (1) found him and killed him, while the fox escaped and was happy to be rid of the greedy wolf.” |
2.2 Generated Text from model mLED, quality level 1-2 |
“The wolf wants to catch the fox because he is the weakest and would be the master. (1) The fox tells the wolf to get a lamb. (1) The lamb is not happy, (3) but the other wants to have it too. The farmers run to get the lamb, and they hit him (2). When they come to rescue them, the witch tells them they are a nimmersatt (3, 4). They go to a farmer’s house, where they (1) cook pancakes, and the sheep eat the pancakes again. (1) The sheep shout loudly and calls for help, but when she comes out, she says that they have been slaughtered by a man in the forest, and he says that he wants to help out. (3) The wolfe lets go and jumps through the hole, but he does not want to see anything. He jumps out of the hole and kills the shepherd, who does not know what to do. (3) On the third day, he (2) is happy that he is rid of the old Nimmersatt (4), who has been rid of a man who is slaughtering meat in the cellar, and who is the first to get rid of him. (3)” |
Example 2 |
@@ 6 Discussion and analysis
6.1 How well do different metrics perform at segment level?
Figure 5 shows Spearman’s ranking correlation between human annotation and evaluation metrics at the segment level. Recall that the correlation results between GPT-4 and human annotation are provided in Table 11 (see Table 21 for GPT-3.5).
Decent correlation from BERTScore variants but worse for German For hDe-En, we observe a weak to moderate level of correlation with human annotation for BERTScore, BARTScore, , and . MoverScore correlates weakly in coherence and relevance and the same is true for ROUGE-L. ROUGE-1 has low correlations with human annotation and the same applies for and DiscoScore. The situation is slightly worse for hEn-De where we observe a lower level of correlation in general compared to hDe-En and less significant correlations from BARTScore and BERTScore variants according to the Student’s t-test especially for coherence. ROUGE-L correlates moderately for direction hEn-De in consistency and relevance. Overall, BERTScore variants correlate slightly better than other metrics with humans in hDe-En for coherence (0.67), consistency (0.61), and relevance (0.66), and in hEn-De for consistency (0.41) and relevance (0.37).
6.2 Statistical analysis on model performance
To quantify the effect of document level characteristics on model performance and contribute to explainability of our models, we analyze how BERTScore-F1 as our best automatic metric overall changes depending on four document features reported in Section 3, namely the mean length of documents (Length), mean dependency distance (MDD), publication year (Year) and divergence of sentence semantics between document and summary by computing the mean sentence-level cosine similarity (Similarity). We apply normalization to numerical features including BERTScore-F1, Length, MDD and Similarity to have zero mean and unit variance.181818We exclude the concern of multicollinearity for numerical variables by computing variance inflation factors (VIF) and the details are reported in Section 9.8 (appendix). We use publication year as a categorical variable where we categorize it by two time periods for hDe-En, namely 1800-1850 (Year:1800-1850) and after 1850 (Year:1850+), and for hEn-De we add another grouping which includes all documents before 1800 (Year:-1800). To examine model effects, we also encode a categorical variable to represent different models (Model). The regression formula is thus:
where represents documents, represents year groups, and represents models defined in Section 4 and
is an error term. We fit Equation (1) using ordinary least squares via statsmodels Seabold and Perktold (2010) separately for both directions (hDe-En and hEn-De).
Direction | Numerical | Categorical | |||
hDe-En | 0.06∗∗∗ | -0.14∗∗∗ | 0.03∗ | base | -0.05 |
hEn-De | 0.24∗∗∗ | -0.15∗∗∗ | -0.08∗∗∗ | -0.003 | 0.69∗∗∗ |
Effect of document features We report the regression results of document features in Table 13.191919The adjusted is 0.425 for direction hDe-En and 0.551 for direction hEn-De. For both directions, embedding similarity measured by has a significantly positive effect on model performance which indicates the model performs better with higher embedding similarity between document and summary. In contrast, the mean length of document (measured by ) shows a significantly negative relation to model performance where more tokens in the source document result in worse performance. This is expected since longer inputs increase the difficulty of summarization and also increase the chance of information loss due to input truncation. The effect of mean dependency distance () varies. We obtain a significantly negative effect for hEn-De which indicates that sentences with more syntactic complexity have a negative influence on model performance. A positive effect for hDe-En is obtained but with a low level of significance. The effect of document publication year is not significant for hDe-En where the model performance differs insignificantly between base group Year:1800-1850 and group Year:1850+. For hEn-De, this also holds for documents published in 1800-1850 compared to the base group with the publication year before 1800. However, we observe a significantly positive effect on model performance if we compare the group Year: 1850+ to base group Year:-1800. This effect also coincides with our finding from Section 3 that English documents from the late 1800s share more similarities to modern English, and thus have a positive impact on model performance.
Also worth noticing is that the main contributor among document features to model performance varies between directions. For hDe-En, the main influence comes from the length of document and mildly from embedding similarity while for hEn-De, the main contributor is the publication year (group Year:1850+) where embedding similarity and the length of document rank the 2nd and 3rd among document features respectively. Overall, however, our findings are largely consistent and they are meaningful: as we argued in the introduction, older texts are more difficult to handle, and the same is true for longer and more syntactically complex source documents; more generally, the more divergence between source documents and reference, the harder is the summarization task.
Effect of model choice
The effect of model choice coincides with our findings from Section 5 where for both directions, we obtain a significant gain with ChatGPT models (except for ChatGPT title prompt) and in contrast, extractive models (extract-then-translate) are significantly worse than the baseline model mLED-Base. For hDe-En, the ChatGPT title prompt model performs significantly worse than the baseline mLED-Base, though not significant for hEn-De. We also notice that for hDe-En, models with intermediate finetuning provide significantly worse performance compared to the baseline as discussed in Section 5 and for hEn-De, the improvement of model performance adding intermediate finetuning tasks is not significant. We include the complete results in Section 9.8 (appendix).
Effect of training set size As discussed in Section 3 and Table 2, the size of our corpora is much smaller compared to other datasets such as CNN/Daily Mail, Wikilingua, and MLSUM. To investigate whether the limited amount of data for finetuning is the main cause of bad performance, we harvest additional instances for both directions. Specifically, we expand the existing datasets for hDe-En from 328 to 455 instances (38.6% increase) and for hEn-De from 289 to 501 instances (73.4% increase). The newly added sources for hDe-En mainly include 95 chapters from the philosophical novel Wilhelm Meister’s Apprenticeship published in 1795, 12 books of epic poetry Reineke Fuchs published in 1794, and 15 scenes from Friedrich Schiller’s play The robbers published in 1781. For hEn-De, the new sources consist of 210 historical translations of fairy tales from different parts of the world published between 1885 and 1910 and 2 short stories.202020The publication time of the newly acquired datasets for both directions closely aligns with that of the original dataset discussed in Section 3. We retrain the CLCTS baseline models using the expanded datasets for both directions, employing the same methods described in Section 4. To ensure comparability, we perform automatic evaluations on the same test sets. Table 19 shows the results for both directions. For hDe-En, the baseline models exhibit marginal improvements, with a mean increase of 0.004 in ROUGE-1, 0.010 in ROUGE-L, and 0.008 in BERTScore-F1. Conversely for hEn-De, the model performances stay invariant on average according to ROUFE-L and BERTScore-F1 with a slight decrease according to ROUGE-1 (-0.004). We speculate that the decrease is because the original CLCTS hEn-De datasets contain more diverse sources (fairy tales, short stories, and plays) while the newly acquired training set primarily contains fairy tales. Moreover, as demonstrated in Section 6.1, the correlation between automatic evaluation metrics and human annotation is worse for hEn-De, which may also affect the evaluation. Overall, our experiments with a larger training set size only show marginal improvements to the existing baseline models. Based on this observation, we hypothesize that a greater volume of data might be necessary to effectively address the intricate nature of the CLCTS task.